OpenStudy (anonymous):

can someone pLease heLp me ! Either in your home or in a store, find a container of salt. Find the mass (or weight) of the salt on the package. Refer to the Periodic Table to get the atomic masses of sodium (Na) and chlorine (Cl). Add them together to get the formula weight of sodium chloride (table salt). Divide the total mass in grams of the package of salt by the formula weight to get the amount of salt in moles. You may have to convert ounces to grams before doing the calculation. (Use 28.4 grams to an ounce as a conversion factor.) Purpose: To convert from grams to moles, find the

OpenStudy (anonymous):

Purpose: To convert from grams to moles, find the formula mass of a compound , and determine the percentage of mass of the elements in a compound. Materials: container of salt (weight 737 grams), Periodic Table of Elements. http://www.ptable.com/ , and a calculator. Tip: When finding the mass of the sodium and chloride ions make sure that you use the atomic mass units: Cl : 35.35 grams and Na: 22.99 grams. This will not be your answer. You are going to use the mass of Cl and Na to get to the total mass of the ions. Procedure: Hypothetical Example: Carbon monoxide, CO (You will use NaCl instead) Question One: How much salt, in moles, is there in the container? Step 1: Find the mass or weight on the container. 50 grams. (Use 737 grams of NaCl)

OpenStudy (anonymous):

Step 2: Use the periodic table to find the atomic masses of each of the elements in the compound. C= 12.01 grams O= 16.00 grams Na = ? grams Cl = ? grams Step 3: Add all of the atomic masses together to get the formula weight. 12.01 grams + 16.00 grams = 28.01 grams ? grams of Na + ? grams of Cl = formula weight Step 4: To get the number of moles (FINAL ANSWER) (mass of the package)/ (formula weight) = # of moles (50 grams)/ (28.01 grams) = 1.8 moles 737 grams / ? grams of NaCl = ? moles NaCl Question Two: How much mass is there of sodium ions? Chloride ions? Step 5: CO has a 1:1 ratio of moles. There are 1.8 moles of Carbon and 1.8 moles of Oxygen Step 6: Determine the mass of carbon ions: (1.8 mol C)x (12.01 grams C)/ (1 mol) = 21.6 grams of C ( ? mol Na) X ( ? grams Na) / (1mol) = ? grams of Na Determine the mass of oxygen ions: (1.8 mol O) x (16.00 grams O)/(1 mol) = 28.8 grams of O ( ? mol Cl) X ( ? grams Cl) / (1mol) = ? grams of Cl

OpenStudy (anonymous):

*Check your answer: take the mass of the carbon ions plus the mass of the oxygen ions to get the mass of the container.* 21.6 grams of C + 28.8 grams of O = 50 grams (the weight in the container) Question Three: What percentage of the total mass is sodium? Chlorine? Step 7: To find the percentage of oxygen and the percentage of carbon use this equation: % mass = (mass of the element ion)/ (weight of the container) x 100 % mass of carbon = (21.6 grams C)/(50 grams)x 100 = 43.2 % C % mass of sodium = ( ? grams Na)/ (737 grams) x 100 =? % Na % mass of oxygen = (28.8 grams O)/(50 grams) x 100 = 57.6 % O % mass of chlorine = ( ? grams Cl)/ (737 grams) x 100 =? % Cl *check your answer % mass should total approximately 100% *